# STG-Mamba: Spatial-Temporal Graph Learning via Selective State Space Model

### v2

@@ Abstract
Spatial-Temporal Graph (STG) data is characterized as dynamic, heterogenous, and non-stationary, leading to the continuous challenge of spatial-temporal graph learning. In the past few years, various GNN-based methods have been proposed to solely focus on mimicking the relationships among node individuals of the STG network, ignoring the significance of modeling the intrinsic features that exist in STG system over time. In contrast, modern Selective State Space Models (SSSMs) present a new approach which treat STG Network as a system, and meticulously explore the STG system’s dynamic state evolution across temporal dimension. In this work, we introduce Spatial-Temporal Graph Mamba (STG-Mamba) as the first exploration of leveraging the powerful selective state space models for STG learning by treating STG Network as a system, and employing the Graph Selective State Space Block (GS3B) to precisely characterize the dynamic evolution of STG networks. STG-Mamba is formulated as an Encoder-Decoder architecture, which takes GS3B as the basic module, for efficient sequential data modeling. Furthermore, to strengthen GNN’s ability of modeling STG data under the setting of SSSMs, we propose Kalman Filtering Graph Neural Networks (KFGN) for adaptive graph structure upgrading. KFGN smoothly fits in the context of selective state space evolution, and at the same time keeps linear complexity. Extensive empirical studies are conducted on three benchmark STG forecasting datasets, demonstrating the performance superiority and computational efficiency of STG-Mamba. It not only surpasses existing state-of-the-art methods in terms of STG forecasting performance, but also effectively alleviate the computational bottleneck of large-scale graph networks in reducing the computational cost of FLOPs and test inference time.
@@ 1 Introduction
Spatial-temporal graph data is a kind of Non-Euclidean data which widely exists in our daily life, such as urban traffic network, metro system InFlow/OutFlow, social networks, regional energy load, weather observations, etc. Owning to the strong ability of capturing dynamic features across both temporal and spatial dimension, attention-based methods, such as Spatial-Temporal Transformer and its variants have recently been broadly implemented in spatial-temporal graph (STG) learning tasks. In the domain of STG traffic forecasting, seminal works have emerged that leverage these methods to enhance prediction accuracy. ASTGCN Guo et al. (2019) introduces an innovative approach combining spatial-attention & temporal attention mechanisms with graph neural networks. ADCT-Net Kong et al. (2024) presents to incorporate a dual-graphic cross-fusion Transformer for dynamic traffic forecasting. Huo et al. (2023) integrates hierarchical graph convolutional networks with Transformer networks to capture sophisticated STG relationships among traffic data. For weather forecasting, Liang et al. (2023) exemplifies how Transformer-based models can be utilized for accurate air quality forecasting across vast geographical areas. Chen et al. (2023) further demonstrates the power of GNNs in predicting air quality, emphasizing the importance of group-aware mechanisms in enhancing STG forecasting performance. In the realm of Social Recommendation, Huang et al. (2021) introduces a coupled GNN framework that incorporates inter-dependent knowledge graphs across both items and users. Wang et al. (2020a) explores the use of disentangled representations within a heterogeneous graph attention network to refine recommendation systems.
@@
Lately, with the popularization of Mamba Gu and Dao (2023); Wang et al. (2024); Liu et al. (2024), Selective State Space Models (SSSM) have aroused considerable interest among researchers from computer vision and Natural Language Processing. SSSM is a variant of the State Space Models, which originated from Control Science and Engineering fields Lee et al. (1994); Friedland (2012) with strong theoretical foundation. State Space Models provide a professional framework to describe a physical system’s dynamics state evolution through sets of input, output, and state variables related by first-order differential or difference equations. The approach allows for a compact way to model and analyze systems with multiple inputs and outputs (MIMO) Aoki (2013).
As a pioneer deep learning-based SSSM, Mamba is proposed to address data-centric context learning for sequential tasks. Different from attention score calculation, Mamba construct selective state space models for the research objects, which encode contexts by using recurrent hidden states. The selection mechanism within Mamba allows it to control which part of the input data should be flowed into hidden states, as the significant context features that will influence the next-step embedding update.
@@
  • We make the first attempt to adapting SSSMs for STG learning tasks. A simple yet elegant way is employed to extend selective state space models to handle STG data. Specifically, we formulate the framework as Encoder-Decoder architecture, with $N$ stacked GS3B as basic modules and the proposed Kalman Filtering Graph Neural Network (KFGN) within it, to facilitate efficient Seq2Seq modeling.
  • A novel Graph Selective State Space Block (GS3B) is proposed for STG network pioneering integration with selective state space models, which performs input-dependent edge construction and dynamic node feature selection. GS3B is capable of capturing the sophisticated heterogenous dependencies among STG network, and accurately modeling the dynamic evolution of STG system over time, making it highly fitted for STG learning.
  • We introduce Kalman Filtering Graph Neural Network (KFGN) as the specialized adaptive spatiotemporal graph generation and upgrading method, which smoothly fits in within the SSSM-based context, ensuring the synchronous updates between graph structure and STG systems’s current state.
@@ 2 Preliminaries
Spatial-Temporal Graph (STG). STG is a kind of non-Euclidean structured data, which extensively existed in daily-life, business, and society. For a given graph network, if the node individuals have certain spatial connectivities at any given time step $\tau$, and the node features and/or graph structures are dynamically evolving over temporal axis, it could be defined as an STG network. Formally, we use $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ to represent a given STG network. Here, $V^{ST}$ denotes the set of graph vertices (nodes) within the network, $E^{ST}$ denotes the edges of the graph network. $A^{ST}$ is the graph adjacency matrix, which indicates the calculated proximity between node individuals.
@@
State Space Models. State Space Models (SSMs) are a foundational concept in control science and system identification, offering a comprehensive framework to describe the dynamic evolutions of systems through mathematical theories. SSMs articulate a system’s behavior by defining a set of equations that represent the system’s state evolution over time, incorporating both control inputs and outputs.
Formally, SSMs are expressed through state equations $x_{t+1}=f(x_{t},u_{t})$ and observation equations $y_{t}=g(x_{t},u_{t})$, where $x_{t}$ denotes the state at time $t$, $u_{t}$ represents the control input, and $y_{t}$ is the observed output. The state transition function $f(\cdot)$ describes how the current state $x_{t}$ evolves to the next state $x_{t+1}$ under the influence of the control input $u_{t}$. $f(\cdot)$ captures the dynamics of the system, accounting for how internal and external factors modify the system’s state over time. The observation function $g(\cdot)$ is used for mapping the current state $x_{t}$ and control input $u_{t}$ to the observed output $y_{t}$. This function defines how the underlying state of the system is reflected in the measurements or observations collected.
As a variant of SSMs, Selective State Space Model emphasizes how to build the selection mechanism upon SSMs, which is highly resemble to the core idea of attention mechanism, making it a competitive counterpart of Transformer architecture.
STG Forecasting based on SSSMs. When employing Selective State Space Models (SSSMs) for spatial-temporal graph forecasting, the problem can be formulated as dynamically identifying and utilizing relevant portions of historical spatiotemporal data and graph structures to predict the future states of a STG system.
@@ 3 Methodology
In this section, we start with introducing the model architecture of STG-Mamba and the formulation of STG Selective State Space Models. Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable dynamic adaptive graph learning; and the Graph Selective State Space Blocks (GS3B) which acts as the basic modules for STG-Mamba (Section 3.3). Finally, we describe the computational efficiency of STG-Mamba in Section 3.4.
@@ 3.1 Formulation of STG Selective State Space Models
Modern State Space Models all rely on a classical continuous system that maps a one-dimensional input function or sequence, denoted as $x(t)\in\mathbb{R}$, through intermediate implicit states $h(t)\in\mathbb{R}^{N}$ to an output $y(t)\in\mathbb{R}$. The aforementioned process can be represented as a linear Ordinary Differential Equation (ODE) as in Eq. 2:
$\begin{array}[]{lr}h^{\prime}(t)=\mathbb{A}h_{t}+\mathbb{B}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (2)
where $\mathbb{A}\in\mathbb{R}^{N\times N}$ represents the State Matrix, while $\mathbb{B}\in\mathbb{R}^{N\times 1},\mathbb{C}\in\mathbb{R}^{N\times 1}$ denote the Input Matrix and Output Matrix, respectively.
Obtaining the output sequence $y(t)$ at a given time-step $t$ necessitates determining $h(t)$ first, which poses great analytical challenges. As such, modern selective state space models address this challenge by converting the continuous system into a discretized version, enhancing the suitability for deep learning scenarios while boosting computational efficiency. A critical step in the discretization process involves the introduction of a time-scale factor $\Delta$, leading to the discrete representation of matrices $\mathbb{A}$ and $\mathbb{B}$ as $\overline{\mathbb{A}}$ and $\overline{\mathbb{B}}$, respectively, following a pre-determined discretization method, commonly the zero-order hold (ZOH) approach.
$\begin{array}[]{lr}\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})\\ \overline{\mathbb{B}}=(\Delta\mathbb{A})^{-1}(\exp(\Delta A)-\mathbb{I})\cdot\Delta\mathbb{B}\end{array}$ (3)
After discretization, SSM-based methods can be computed in two ways: linear recurrent or global convolution, defined as in Eq. 4 and Eq. 5, respectively:
$\begin{array}[]{lr}h^{\prime}(t)=\overline{\mathbb{A}}h_{t}+\overline{\mathbb{B}}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (4)
$\begin{array}[]{lr}\overline{\mathbb{K}}=(\mathbb{C\overline{B}},\mathbb{C\overline{AB}},...,\mathbb{C\overline{A}^{L-1}\overline{B}})\\ y=x*\overline{\mathbb{K}}\end{array}$ (5)
where $\overline{K}\in\mathbb{R}^{L}$ denotes a structured convolutional kernel, and $L$ denotes the length of the input sequence $x$.
Figure 1 illustrates the comprehensive architecture of the proposed STG-Mamba. Specifically, we formulate the overall architecture following the Encoder-Decoder fashion for efficient sequential data modeling and prediction. In STG-Mamba, we leverage the Graph Selective State Space Block (GS3B) as the basic module, and repeat it $N$ times for Encoder and Decoder part. GS3B consists of several operations, including Layer Norm, Linear Layer, Kalman Filter Graph Neural Networks (KFGN), Graph State Space Layer (GSS-Layer), and Residual Connections. We specifically design KFGN and GSS-Layer for STG learning task build upon selective state space models.
@@ 3.2 Kalman Filtering Graph Neural Networks
Due to the dynamic, heterogenous, and non-stationary characteristics of STG data, the spatiotemporal dependencies of an STG system also evolve over time. When reflecting such graph dynamics in GNN formulation, it means the time-variant dynamic graph adjacency matrix. In existing literature, researchers usually employ various kinds of Attention-based methods to construct dynamic graph adjacency matrices, which may not be feasible in the context of state space model-based STG learning. On the other hand, Kalman Filtering (KF), as a well-known application of SSMs, has been extensively used in sequential data filtering and forecasting.
In light of that, we propose a novel approach which employs KF-based method to generate dynamic-evolving graphs, which delicately captures the system’s state-space transition for a given STG network. The method is termed as Kalman Filtering Graph Neural Networks (KFGN). To generate graph adjacency matrix $\mathbb{\alpha_{t}}$ at each time step using KFGN, the modeling procedure involves two steps as shown in the following:
Step 1: State Equation Calculation: Let $\mathbb{\alpha_{t}}$ denotes the graph adjacency matrix at time $t$, which indicates the current state of the STG system. The state equation is defined as $\mathbb{\alpha_{t}}=F_{t-1}\mathbb{\alpha_{t-1}}+W_{t-1}$, where $F_{t-1}$ is the state transition matrix that reflects changes in the graph adjacency matrix from $t-1$ to $t$, and $W_{t-1}$ represents the noise of the KF transition procedure.
Step 2: Observation Equation Calculation: The observed STG feature can be modeled as $X_{t}^{{}^{\prime}}=H_{t}\mathbb{\alpha_{t}}+V_{t}$, where $X_{t}^{{}^{\prime}}$ is the observation of STG feature at time $t$, $H_{t}$ is the observation matrix, and $V_{t}$ represents observation noise.
Iteratively applying these two equations through Kalman Filter enables dynamic estimation of the graph adjacency matrix $\alpha_{t}$ at each moment, adaptively capturing the evolving dynamics of the STG network. In practice, we employ a Neural Network-based method called Deep Kalman Filter (DKF) to automatically learn the state transition matrix $F_{t-1}$ and observation matrix $H_{t}$. Specifically, we adopt Gated Recurrent Unit (GRU) network as the architecture of Deep Kalman Filter. The GRU-based DKF is employed to learn the non-linear mapping for state transition and observation generation. The state transition matrix $F_{t-1}$ is modeled by GRU inner process to capture the non-linear relationships between states, predicting the state for the next moment. Similarly, the observation matrix $H_{t}$ is predicted under the current state by GRU. This process enables the system to automatically learn the dynamics of state transitions and observation generation from data without manually settings.
After receiving the $\alpha_{t}$, the next step of KFGN is the graph convolution operation, which is defined as:
$h_{t}^{kfgn}=(\omega_{gc}\odot\alpha_{t})x_{t}^{{}^{\prime}}$ (6)
where $\odot$ is the Hadamard product, and $x_{t}^{{}^{\prime}}$ denotes the input feature of KFGN at time $t$. $\omega_{gc}$ is the learnable weight matrix of KFGN.
3.3 GS3B: Graph Selective State Space Blocks

### v3

@@ Abstract
Spatial-Temporal Graph (STG) data is characterized as dynamic, heterogenous, and non-stationary, leading to the continuous challenge of spatial-temporal graph learning. In the past few years, various GNN-based methods have been proposed to solely focus on mimicking the relationships among node individuals of the STG network, ignoring the significance of modeling the intrinsic features that exist in STG system over time. In contrast, modern Selective State Space Models (SSSMs) present a new approach which treat STG Network as a system, and meticulously explore the STG system’s dynamic state evolution across temporal dimension. In this work, we introduce Spatial-Temporal Graph Mamba (STG-Mamba) as the first exploration of leveraging the powerful selective state space models for STG learning by treating STG Network as a system, and employing the Spatial-Temporal Selective State Space Module (ST-S3M) to precisely focus on the selected STG latent features. Furthermore, to strengthen GNN’s ability of modeling STG data under the setting of selective state space models, we propose Kalman Filtering Graph Neural Networks (KFGN) for dynamically integrate and upgrade the STG embeddings from different temporal granularities through a learnable Kalman Filtering statistical theory-based approach. Extensive empirical studies are conducted on three benchmark STG forecasting datasets, demonstrating the performance superiority and computational efficiency of STG-Mamba. It not only surpasses existing state-of-the-art methods in terms of STG forecasting performance, but also effectively alleviate the computational bottleneck of large-scale graph networks in reducing the computational cost of FLOPs and test inference time. The implementation code is available at: https://github.com/LincanLi98/STG-Mamba.
@@ 1 Introduction
Spatial-temporal graph data is a kind of Non-Euclidean data which widely exists in our daily life, such as urban traffic network, metro system InFlow/OutFlow, social networks, regional energy load, weather observations, etc. Owning to the strong ability of capturing dynamic features across both temporal and spatial dimension, attention-based methods, such as Spatial-Temporal Transformer and its variants have recently been broadly implemented in spatial-temporal graph (STG) learning tasks. In the domain of STG traffic forecasting, seminal works have emerged that leverage these methods to enhance prediction accuracy. ASTGCN Guo et al. (2019) introduces an innovative approach combining spatial-attention & temporal attention mechanisms with graph neural networks. ADCT-Net Kong et al. (2024) presents to incorporate a dual-graphic cross-fusion Transformer for dynamic traffic forecasting. Huo et al. (2023) integrates hierarchical graph convolutional networks with Transformer networks to capture sophisticated STG relationships among traffic data. For weather forecasting, Liang et al. (2023) exemplifies how Transformer-based models can be utilized for accurate air quality forecasting across vast geographical areas. Chen et al. (2023) further demonstrates the power of GNNs in predicting air quality, emphasizing the importance of group-aware mechanisms in enhancing STG forecasting performance. In the realm of Social Recommendation, Zhang et al. (2023) introduces an Adaptive Disentangled Transformer (ADT) framework that optimizes the disentanglement of attention heads in different layers of Transformer architecture. Wei et al. (2023) explores the improved efficiency of multimedia recommendation by introducing modal-specific embeddings and a lightweight self-attention.
@@
Lately, with the popularization of Mamba Gu and Dao (2023); Wang et al. (2024); Liu et al. (2024), Selective State Space Models (SSSM) have aroused considerable interest among researchers from computer vision and Natural Language Processing. SSSM is a variant of the State Space Models, which originated from Control Science and Engineering fields Lee et al. (1994); Friedland (2012). State Space Models provide a professional framework to describe a physical system’s dynamics state evolution through sets of input, output, and state variables related by first-order differential or difference equations. The approach allows for a compact way to model and analyze systems with multiple inputs and outputs (MIMO) Aoki (2013).
@@
  • We make the first attempt to adapting SSSMs for STG learning tasks. A simple yet elegant way is employed to extend selective state space models to handle STG data. Specifically, we formulate the framework following the Stacked Residual Encoder Fashion, with $N$ stacked Graph Selective State Space Block (GS3B) as basic modules and the proposed Kalman Filtering Graph Neural Network (KFGN) within it, to facilitate efficient Seq2Seq modeling.
  • A novel Spatial-Temporal Selective State Space Module (ST-S3M) is proposed for STG network pioneering integration with selective state space models, which performs input-dependent adaptive spatial-temporal graph feature selection. ST-S3M is capable of capturing the sophisticated heterogenous dependencies among STG network, and accurately modeling the dynamic evolution of STG system over time, making it highly fitted for STG learning.
  • We introduce KFGN as the specialized adaptive spatiotemporal graph generation and upgrading method, which smoothly fits in within the SSSM-based context. In KFGN, the Dynamic Filter Linear serves as dynamic GNN module which generating input-specific graph structures. The KF-Upgrading mechanism models STG inputs from different temporal granularities as parallel streams, and the output embeddings are integrated through the statistical theory learning of Kalman Filtering for optimization.
@@ 2 Preliminaries
Spatial-Temporal Graph (STG). STG is a kind of non-Euclidean structured data, which extensively existed in daily-life, business, and society. For a given graph network, if the node individuals have certain spatial connectivity at any given time step $\tau$, and the node features and/or graph structures are dynamically evolving over temporal axis, it could be defined as an STG network. Formally, we use $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ to represent a given STG network. Here, $V^{ST}$ denotes the set of graph vertices (nodes) within the network, $E^{ST}$ denotes the edges of the graph network. $A^{ST}$ is the graph adjacency matrix, which indicates the calculated proximity between node individuals.
@@
State Space Models (SSMs). Deep Learning-based SSMs are a lately invented class of sequential models that are closely related to RNN architectures and classical state space models. They are characterized by a particular continuous system model that maps a multi-dimensional input sequence to a corresponding output sequence through an implicit latent state representation. SSMs are defined by four parameters $(\mathbb{A},\mathbb{B},\mathbb{C},\mathbb{D})$ that specify how the input (control signal) and the current state determine the next state and output. This framework allows for efficient sequential modeling by enabling both linear and non-linear computations. As a variant of SSMs, SSSMs emphasize how to build the selection mechanism upon SSMs, which is highly resemble to the core idea of attention mechanism, making it a competitive counterpart of Transformer architecture.
STG Forecasting based on SSSMs. Employing SSSMs for spatial-temporal graph forecasting, the problem can be formulated as dynamically identifying and utilizing relevant portions of historical spatiotemporal data and graph structures to predict the future states of a STG system.
@@ 3 Methodology
In this section, we start with introducing the model architecture of STG-Mamba and the formulation of STG Selective State Space Models. Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable adaptive statistical-based graph learning under contexts with noisy data and uncertainty; and the Spatial-Temporal Selective State Space Module (ST-S3M) which serves as the state space selection mechanism for spatial-temporal graph learning. Finally, we describe the computational efficiency of STG-Mamba in Section 3.4.
@@ 3.1 Formulation of STG Selective State Space Models
Modern selective state space models, such as Mamba, build upon classical continuous systems by incorporating deep learning techniques and discretization methods suitable for complex time-series data. Mamba specifically maps a one-dimensional input sequence $x(t)\in\mathbb{R}$ through intermediate implicit states $h(t)\in\mathbb{R}^{N}$, culminating in an output $y(t)\in\mathbb{R}$. The fundamental process is described by a modified set of state space equations that include discretization adjustments to accommodate deep learning frameworks:
$\begin{array}[]{lr}h^{\prime}(t)=\overline{\mathbb{A}}h(t)+\overline{\mathbb{B}}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (2)
where $\overline{\mathbb{A}}\in\mathbb{R}^{N\times N}$ and $\overline{\mathbb{B}}\in\mathbb{R}^{N\times 1}$ are discretized representations of the continuous state transition and input matrices, respectively. $\mathbb{C}\in\mathbb{R}^{1\times N}$ is the output matrices. The discretization process involves transformation techniques that convert $\mathbb{A}$ and $\mathbb{B}$ from their continuous forms to computational efficient formats. These transformations are determined by a selective scan algorithm that utilizes a convolutional approach to model interactions across temporal dimension:
$\begin{array}[]{lr}\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})\\ \overline{\mathbb{B}}=\Delta\mathbb{B}\cdot(\exp(\Delta\mathbb{A})-\mathbb{I})\end{array}$ (3)
Mamba extends classic state space model functionalities by incorporating a selective mechanism that dynamically adapts to the input sequence characteristics. This adaptability is achieved through a novel architecture that selectively applies convolution kernels across the sequence, optimizing the model’s response to spatiotemporal variations:
$\begin{array}[]{lr}\overline{\mathbb{K}}=(\mathbb{C}\overline{\mathbb{B}},\mathbb{C}\overline{\mathbb{A}\mathbb{B}},\dots,\mathbb{C}\overline{\mathbb{A}^{L-1}\mathbb{B}})\end{array}$ (4)
where $\overline{\mathbb{K}}$ denotes a structured convolutional kernel, capturing dependencies across spatial and temporal dimensions, and $L$ represents the length of the input sequence $x$. This selective convolutional approach significantly enhances the model’s capability to forecast spatiotemporal data by focusing computation on critical features of the data sequence, thus improving both accuracy and efficiency in predictions.
Figure 1 illustrates the comprehensive architecture of the proposed STG-Mamba. Specifically, we formulate the overall architecture following the Residual Encoder fashion for efficient sequential data modeling and prediction. In STG-Mamba, we leverage the Graph Selective State Space Block (GS3B) as the basic Encoder module, and repeat it $N$ times. GS3B consists of several networks and operations, including Layer Norm, Kalman Filter Graph Neural Networks (KFGN) which consists of DynamicFilter-GNN followed by Kalman Filter Upgrading (KF-Upgrading), Spatial-Temporal Selective State Space Module (ST-S3M) which consists of Linear & Split, Conv1D, SiLU, State Space Selection Mechanism (SSSM), Element-Wise Concatenation, as well as th Residual Connections for every GS3B Encoder layer.
@@ 3.2 Kalman Filtering Graph Neural Networks
The motivation for employing Kalman Filtering-based methods in constructing GNNs stems from the need to enhance the reliability and accuracy of spatial-temporal forecasting. STG big data (such as traffic sensor records/weather station records/urban metro system In/Out records) often contains inherent biases and noises that existing methods typically overlook. By integrating Kalman Filtering-based optimization and upgrading, our KFGN effectively addresses these inaccuracies by dynamically weighing the reliability of data streams from different temporal granularity, optimizing the fusion of these data streams based on the estimated variances. This approach not only corrects dataset inherent errors but also significantly improves the model’s capacity to capture the complex inter-dependencies within STG patterns.
As depicted in Figure 2, the KFGN pipeline consists of two key steps. In the first step, embeddings generated from the model inputs of different temporal granularities (i.e., recent steps/periodic steps/trend steps) are sent into a DynamicFilter-GNN module for processing. Then, in the second step, the outputs of DynamicFilter-GNN module are integrated and optimized through the Kalman Filtering Upgrading module. In the initial stage of DynamicFilter-GNN, a learnable Base Filter parameter matrix of size $\mathbb{R}^{\text{in\_fea}\times\text{in\_fea}}$ is defined. It’s employed to transform the graph adjacency matrix, thereby dynamically adjusting the connection degrees between nodes. Following that, we uniformly initialize the weights and biases, with an initialization standard deviation stdv calculated as the reciprocal of the square root of the number of input features: $\text{stdv}=\frac{1}{\sqrt{\text{in\_fea}}}$. This uniform initialization is critical for ensuring the module starts from a neutral point.
Figure 2: Kalman Filtering Graph Neural Networks.
Subsequently, we transformer the Base Filter via a linear transformation layer, and combine it with the original adjacency matrix $A_{\text{ini}}^{c_{i}}$, resulting in a dynamically adjusted adjacency matrix $A_{\text{dyn}}^{c_{i}}$. Let $c_{i,i=\{1,2,3\}}\in\{r,p,q\}$ represents the input data from different temporal granularities, (i.e., $c_{1}=r$ denotes historical recent, $c_{2}=p$ denotes historical period, $c_{3}=q$ denotes historical trend). Leveraging $A_{\text{dyn}}^{c_{i}}$, input embedding $h_{in}^{c_{i}}$, weights $W_{DF}^{c_{i}}$ and bias $b_{DF}^{c_{i}}$, the input embedding undergoes a graph convolution:
$h_{DF}^{c_{i}}=h_{in}^{c_{i}}\cdot(A_{\text{dyn}}^{c_{i}}\cdot W_{DF}^{c_{i}})+b_{DF}^{c_{i}}$ (5)
This design enables the model to dynamically adjust the strengths of connections between nodes within the graph based on the learned adjustments.
Following DynamicFilter-GNN, the next step is Kalman Filtering Upgrading. Since the input data are subjected to the same dataset but within different temporal granularities, they are assumed to follow Gaussian Distributions:
$y_{q}(x;\mu_{q},\sigma_{q})\triangleq\frac{\exp^{-\frac{(x-\mu_{q})^{2}}{2\sigma_{q}^{2}}}}{\sqrt{2\pi\sigma_{q}^{2}}};y_{p}(x;\mu_{p},\sigma_{p})\triangleq\frac{\exp^{-\frac{(x-\mu_{p})^{2}}{2\sigma_{p}^{2}}}}{\sqrt{2\pi\sigma_{p}^{2}}};y_{r}(x;\mu_{r},\sigma_{r})\triangleq\frac{\exp^{-\frac{(x-\mu_{r})^{2}}{2\sigma_{r}^{2}}}}{\sqrt{2\pi\sigma_{r}^{2}}};$ (6)
where subscript $q=c_{3},p=c_{2},r=c_{1}$ denote historical trend/period/recent dataset, respectively.
We employ Kalman Filtering method to derive accurate information from the multi-granularity observation sets. Specifically, their individual probability distribution functions can be integrated together by multiplication:
$y_{fuse}(x;\mu_{q},\sigma_{q},\mu_{p},\sigma_{p},\mu_{r},\sigma_{r})$ $=\frac{\exp^{-\frac{(x-\mu_{q})^{2}}{2\sigma_{q}^{2}}}}{\sqrt{2\pi\sigma_{q}^{2}}}\times\frac{\exp^{-\frac{(x-\mu_{p})^{2}}{2\sigma_{p}^{2}}}}{\sqrt{2\pi\sigma_{p}^{2}}}\times\frac{\exp^{-\frac{(x-\mu_{r})^{2}}{2\sigma_{r}^{2}}}}{\sqrt{2\pi\sigma_{r}^{2}}}$
$=\frac{1}{(2\pi)^{3/2}\sqrt{\sigma_{q}^{2}\sigma_{p}^{2}\sigma_{r}^{2}}}\exp^{-\left(\frac{(x-\mu_{q})^{2}}{2\sigma_{q}^{2}}+\frac{(x-\mu_{p})^{2}}{2\sigma_{p}^{2}}+\frac{(x-\mu_{r})^{2}}{2\sigma_{r}^{2}}\right)}$ (7)
By reorganizing Eq. 7 into a simplified version, we have:
$y_{fuse}(x;\mu_{fuse},\sigma_{fuse})=\frac{1}{\sqrt{2\pi\sigma_{fuse}^{2}}}\exp^{-\frac{(x-\mu_{fuse})^{2}}{2\sigma_{fuse}^{2}}}$ (8)
where $\mu_{fuse}=\frac{\mu_{q}/\sigma_{q}^{2}+\mu_{p}/\sigma_{p}^{2}+\mu_{r}/\sigma_{r}^{2}}{1/\sigma_{q}^{2}+1/\sigma_{p}^{2}+1/\sigma_{r}^{2}}$ and $\sigma_{fuse}^{2}=\frac{1}{1/\sigma_{q}^{2}+1/\sigma_{p}^{2}+1/\sigma_{r}^{2}}$.
To simplify the representation of $\mu_{fuse}$ and $\sigma_{fuse}^{2}$, we hereby introduce parameters $\omega_{q}=\frac{1}{\sigma_{q}^{2}},\omega_{p}=\frac{1}{\sigma_{p}^{2}},\omega_{r}=\frac{1}{\sigma_{r}^{2}}$. Then, $\mu_{fuse}$ and $\sigma_{fuse}^{2}$ can be re-written as:
$\mu_{fuse}=\frac{\mu_{q}\omega_{q}+\mu_{p}\omega_{p}+\mu_{r}\omega_{r}}{\omega_{q}+\omega_{p}+\omega_{r}};\\ \sigma_{fuse}^{2}=\frac{1}{\omega_{q}+\omega_{p}+\omega_{r}}$ (9)
This means that observations from different branches can be effectively integrated using weighted sum, and the weight is a combination of variances:
$\mu_{fuse}$ $=$ $\mu_{q}(\frac{\omega_{q}}{\omega_{q}+\omega_{p}+\omega_{r}})+\mu_{p}(\frac{\omega_{p}}{\omega_{q}+\omega_{p}+\omega_{r}})+\mu_{r}(\frac{\omega_{r}}{\omega_{q}+\omega_{p}+\omega_{r}})$ (10)
$\Downarrow$
$y_{fuse}$ $=$ $y_{q}(\frac{\omega_{q}}{\omega_{q}+\omega_{p}+\omega_{r}})+y_{p}(\frac{\omega_{p}}{\omega_{q}+\omega_{p}+\omega_{r}})+y_{r}(\frac{\omega_{r}}{\omega_{q}+\omega_{p}+\omega_{r}})$
Directly calculating the variance of an observation set requires the complete data, which is computationally expensive. Therefore, we estimate the variance by computing the variance distribution of each training samples:
$\mathbb{E}[\sigma_{\{q,p,r\}}^{2}]=\frac{1}{N_{m}}\sum_{i}\frac{(S_{i}-\overline{S})^{2}}{L}$ (11)
where $L$ is the length of each sample sequence, and $N_{m}$ is the number of data samples. $S_{i}$ denotes the $i$-th observation, and $\overline{S}$ denotes the average value of all observed samples. To further improve the integration and upgrading, we add two learnable weight params $\epsilon,\varphi$ to balance different observation branches. To this end, based on the formulation of $y_{fuse}$ in Eq. 10, the output of Kalman Filtering Upgrading module is:
$\tilde{y}_{fuse}=\frac{\epsilon\cdot(\hat{y}_{q}\omega_{q})+\varphi\cdot(\hat{y}_{p}\omega_{p})+\hat{y}_{r}\omega_{r}}{\omega_{q}+\omega_{p}+\omega_{r}}$ (12)
where $\omega_{q}=\frac{1}{\sigma_{q}^{2}},\omega_{p}=\frac{1}{\sigma_{p}^{2}},\omega_{r}=\frac{1}{\sigma_{r}^{2}}$, $\epsilon$ and $\varphi$ are trainable weights. Finally, for the convenience of neural network training and ensure the scalability, we remove the constant denominator $(\omega_{q}+\omega_{p}+\omega_{r})$ in Eq. 12. As such, the final output of Kalman Filtering Upgrading is defined as:
$\hat{y}_{fuse}=\epsilon\cdot(\hat{y}_{q}\omega_{q})+\varphi\cdot(\hat{y}_{p}\omega_{p})+\hat{y}_{r}\omega_{r}$ (13)
Here, $\hat{y}_{fuse}$ is the final output of KF-Upgrading module. Note that our approach is an optimized version of the classic Kalman Filtering theory, which is specifically designed for deep learning-based methodology, ensuring both computationally efficiency, dynamic hierarchical STG feature fusion, and improved accuracy.
3.3 ST-S3M: Spatial-Temporal Selective State Space Module